OpenStudy (anonymous):

The degree measure of each of the three angles of a triangle is an integer. Which of the following cannot be the ratio of their measures? A) 2:3:4 B)3:4:5 C)4:5:6 D)5:6:7 E)6:7:8

OpenStudy (dumbcow):

E) because 6x+7x+8x =180 21x=180 x=180/21 which is not an integer thus 6x = 6*(180/21) is not an integer